# Illuminating Scalar Dark Matter Co-Scattering in EFT with Monophoton Signatures

### v1

@@ 4 Collider Constraints
Following the selection cuts used in the ATLAS search ATLAS:2020uiq , we calculated the selection efficiencies for our signal model. We demand that the leading photon has $E_{T}^{\gamma}>150$ GeV, $1.52<|\eta^{\gamma}|<2.37$ or $|\eta^{\gamma}|<1.37$, and $\Delta\phi(\gamma,E_{T}^{\rm miss})>0.4$. Then we calculate the signal events in the seven SRs and set limit in the $M_{N_{2}}$ vs $(c_{3}^{\prime})_{UL}$ plane as shown in Fig. 9, where $(c_{3}^{\prime})_{UL}$ represent the $95\%$ CL upper limit on $c_{3}^{\prime}$. Here we only show the strongest and weakest limit. We also calculate the HL-LHC projection, for this we scale the observed upper limit on the visible cross section by a factor $\sqrt{139/3000}$ and compare with the theory prediction. Fig. 9 is for DM mass $M_{\chi}=100$ GeV, $M_{N_{1}}=105$ GeV and $Y_{22}^{\prime}=10^{-9}\leavevmode\nobreak\ \rm{GeV}^{-1}$. Here the solid lines indicate current LHC limit while dashed lines are for HL-LHC projection. The strongest limit is set in SRI4 (black) and the weakest limit in SRE1 (red). $(c_{3}^{\prime})_{UL}$ is determined by the production cross section $\sigma(pp\to N_{1}N_{2})$, $Br(N_{2}\to N_{1}\gamma)$ and cut efficiency $\epsilon$. For a fixed $M_{N_{1}}$ and $c_{3}^{\prime}$, the $\sigma(pp\to N_{1}N_{2})$ drops with $M_{N_{2}}$. While, $Br(N_{2}\to N_{1}\gamma)$ remains steady except near the kinematic threshold, as evident from Fig.6. The cut efficiency is different for different SRs. For SRI4, $\epsilon$ increases with $M_{N_{2}}$ when $M_{N_{1}}$ is fixed and it can vary from $10^{-4}$ for $M_{N_{2}}=110$ GeV, to $0.65$ for $M_{N_{2}}=1000$ GeV. Whereas, for SRE1, $\epsilon$ increases initially with $M_{N_{2}}$ and at a certain point it starts dropping. Therefor, $(c_{3}^{\prime})_{UL}$ varies differently for different SRs. The limit is weakest when $M_{N_{2}}\simeq M_{N_{1}}$ as the photon becomes very soft. For $M_{N_{2}}\simeq 200$ GeV limit on $c_{3}^{\prime}$ becomes strongest. In the region $M_{N_{2}}>200$ GeV, for SRI4, the rise in $\epsilon$ gets canceled by the drop in $\sigma$ leading to a almost flat limit on $c_{3}^{\prime}$ w.r.t $M_{N_{2}}$. Whereas for SRE1, the limit weakens as $\sigma$ drops faster compared to the variation in $\epsilon$. Fig. 9 shows the limits in the case $M_{\chi}=300$ GeV, $M_{N_{1}}=315$ GeV and $Y_{22}^{\prime}=10^{-9}\leavevmode\nobreak\ \rm{GeV}^{-1}$ . The dotted line represents $\Gamma(N_{2})/M_{N_{2}}=0.01$. Finally, in Fig. 10, we show the limit on $c_{3}^{\prime}$ in the $\delta_{1}$ vs $\delta_{2}$ plane, as defined in Eq. 11. Here we consider only SRI4 since it sets the strongest limit.
@@ 5 Numerical Analysis
$\delta_{1}(M_{N_{1}}\leavevmode\nobreak\ \rm{[GeV]})$  $10^{-3}$ , $0.5$] $([\approx M_{\chi},750])$
$\delta_{2}(M_{N_{2}}\leavevmode\nobreak\ \rm{[GeV]})$  $0^{-3} , 11$] $([\approx M_{N_{1}},1500])$
@@
In Fig. 13, we show the allowed range of $M_{N_{1}}$ vs $M_{N_{2}}$ with $c^{\prime}_{3}$ in the color pallet after imposing the relic density constraint as given by Eq. 14. We then impose the ATLAS mono-$\gamma$ limit as discussed in Sec. 4. Figure  13 shows that the mono-$\gamma$ search strongly constrain scenarios with large mass splitting between $M_{N_{1}}$ and $M_{N_{2}}$, these points are shown in red in the figure. Furthermore, as discussed above, the relic density constraint requires at least $N_{1}$ to be present in the early Universe, so that $\chi$ can be diluted through co-annihilation or inverse decay. Thus the mass difference between $M_{N_{1}}$ and $M_{\chi}$ cannot be too large, and we obtain an upper limit on $M_{N_{1}}<600\,\rm{GeV}$. It is important to notice that $N_{2}$ can also participate in DM $\chi$ dilution through co-annihilation or inverse decay. However since the contribution of $N_{1}$ is always present, the upper bound on $M_{N_{2}}$ is less stringent, specifically, $M_{N_{2}}\leq 1000\,\mathrm{GeV}$.
@@
In Fig. 13, we depict the allowed points after the relic density constraint and collider limit in the $\Omega_{\chi}h^{2}$ and $c^{\prime}_{3}$ plane with the color pallet corresponding to the DM mass $M_{\chi}$. The points overlapping with the red line satisfy the observed DM relic density within the $2\sigma$ range. As $c^{\prime}_{3}$ decreases, $N_{1,2}$ undergoes early chemical decoupling from the thermal bath, which in turn does not contribute in $\chi$ dilution neither through co-annihilation nor through co-scattering. We can see that for the range of $M_{\chi}$ that survived $c^{\prime}_{3}\leq 3\times 10^{-4}$ , i.e. $M_{\chi}>200\,\rm{GeV}$, corresponds to $M_{N_{1,2}}>M_{\chi}$. It happens because, for $M_{N_{1,2}}>200\,\rm{GeV}$ additional channels such as $N_{1}N_{2}\to t\bar{t}$ become kinematically accessible. In Fig. 13, we show the variation of the allowed points in $Y^{\prime}_{11}$ - $Y^{\prime}_{22}$ plane with the color pallet corresponding to $\Delta^{\chi}_{1}$. As evident from Fig. 13, for large $Y^{\prime}_{11,22}$, $\chi$ thermalizes with the thermal bath due to rapid decay and inverse decay as $\Delta^{\chi}_{1}\to 0$, and the relic density is governed by the thermal decoupling of $N_{1,2}$. It is the region of parameter space where a co-annihilation process governs $\Omega_{\chi}h^{2}$. On the contrary, for $Y^{\prime}_{11,22}<10^{-8}\,\rm{GeV^{-1}}$, $\chi$ undergoes early chemical decoupling and $\Omega_{\chi}h^{2}$ gets frozen when $\chi$ inverse decay to $N_{1,2}$ ceases. Consequently, in this region of parameter space, the value of $\Delta^{\chi}_{1}>0.5$, indicating that the DM relic density is predominantly determined by the co-scattering mechanism. In Fig 13, we show the discovery prospects via the mono-$\gamma$ signal at the HL-LHC. The dotted-color points in the $M_{N_{1}}$ and $M_{N_{2}}$ plane can be probed in HL-LHC while the red-crossed points remain out of HL-LHC reach. These points correspond to lower values of $M_{N_{2}}-M_{N_{1}}$ and/or lower value of $c^{\prime}_{3}$, which leads to suppressed signal efficiency and/or suppressed production cross section for the signal $pp\to N_{2}N_{2}\to\gamma+E_{T}^{\rm miss}$.
@@ 6 Summary and Conclusion
In this work, we investigate the dark matter production mechanism in an EFT framework where we extent the SM with two singlet fermions $N_{1,2}$ and a real singlet scalar $\chi$, which are $Z_{2}$ odd particles. The singlet scalar $\chi$ serves as a dark matter candidate. Both $\chi$ and $N_{1,2}$ are thermal particles and their interactions in the early Universe controls the DM abundance. We emphasize the role of dimension-5 operators in the co-scattering mechanism. The key parameters on which DM dynamics depend are the masses $M_{N_{1,2}},M_{\chi}$ and couplings $\lambda,c_{3}^{\prime},Y_{11,22}$. Note that, we have assumed $c_{1,2}^{\prime}=0$ throughout our analysis.
@@
As concerns the collider constraints, we find that the dimension-5 operators lead to sizable production of $N_{1,2}$ at the LHC and that the most stringent constraint is coming from the mono-$\gamma$ signature via $pp\to N_{1}N_{2}(\to N_{1}\gamma)\to\gamma+E_{T}^{\rm miss}$. Following the ATLAS analysis, we calculate the sensitivity of this signal which primarily depends on the mass difference $M_{N_{2}}-M_{N_{1}}=\delta_{2}$ (or the photon energy) and $c_{3}^{\prime}$. We calculate the mono-$\gamma$ limit on $c_{3}^{\prime}$ in the $\delta_{2}-\delta_{1}$ plane, where $\delta_{1}$ dictates the relative mass splitting between $M_{N_{1}}$ and $M_{\chi}$. The stronger bound on $c_{3}^{\prime}$ is found for larger $\delta_{2}$ since the energy of the photon grows with $\delta_{2}$.

### v2

@@ 4 Collider Constraints
Following the selection cuts used in the ATLAS search ATLAS:2020uiq , we calculated the selection efficiencies for our signal model. We demand that the leading photon has $E_{T}^{\gamma}>150$ GeV, $1.52<|\eta^{\gamma}|<2.37$ or $|\eta^{\gamma}|<1.37$, and $\Delta\phi(\gamma,E_{T}^{\rm miss})>0.4$. Then we calculate the signal events in the seven SRs and set limit in the $M_{N_{2}}$ vs $(c_{3}^{\prime})_{UL}$ plane as shown in Fig. 9, where $(c_{3}^{\prime})_{UL}$ represent the $95\%$ CL upper limit on $c_{3}^{\prime}$. Here we only show the strongest and weakest limit. We also calculate the HL-LHC projection, for this we scale the observed upper limit on the visible cross section by a factor $\sqrt{139/3000}$ and compare with the theory prediction. Fig. 9 is for DM mass $M_{\chi}=100$ GeV, $M_{N_{1}}=105$ GeV and $y_{22}^{\prime}=10^{-9}\penalty 10000\ \rm{GeV}^{-1}$. Here the solid lines indicate current LHC limit while dashed lines are for HL-LHC projection. The strongest limit is set in SRI4 (black) and the weakest limit in SRE1 (red). $(c_{3}^{\prime})_{UL}$ is determined by the production cross section $\sigma(pp\to N_{1}N_{2})$, $Br(N_{2}\to N_{1}\gamma)$ and cut efficiency $\epsilon$. For a fixed $M_{N_{1}}$ and $c_{3}^{\prime}$, the $\sigma(pp\to N_{1}N_{2})$ drops with $M_{N_{2}}$. While, $Br(N_{2}\to N_{1}\gamma)$ remains steady except near the kinematic threshold, as evident from Fig.6. The cut efficiency is different for different SRs. For SRI4, $\epsilon$ increases with $M_{N_{2}}$ when $M_{N_{1}}$ is fixed and it can vary from $10^{-4}$ for $M_{N_{2}}=110$ GeV, to $0.65$ for $M_{N_{2}}=1000$ GeV. Whereas, for SRE1, $\epsilon$ increases initially with $M_{N_{2}}$ and at a certain point it starts dropping. Therefore, $(c_{3}^{\prime})_{UL}$ varies differently for different SRs. The limit is weakest when $M_{N_{2}}\simeq M_{N_{1}}$ as the photon becomes very soft. For $M_{N_{2}}\simeq 200$ GeV limit on $c_{3}^{\prime}$ becomes strongest. In the region $M_{N_{2}}>200$ GeV, for SRI4, the rise in $\epsilon$ gets canceled by the drop in $\sigma$ leading to a almost flat limit on $c_{3}^{\prime}$ w.r.t $M_{N_{2}}$. Whereas for SRE1, the limit weakens as $\sigma$ drops faster compared to the variation in $\epsilon$. Fig. 9 shows the limits in the case $M_{\chi}=300$ GeV, $M_{N_{1}}=315$ GeV and $y_{22}^{\prime}=10^{-9}\penalty 10000\ \rm{GeV}^{-1}$ . The dotted line represents $\Gamma(N_{2})/M_{N_{2}}=0.01$. Finally, in Fig. 10, we show the limit on $c_{3}^{\prime}$ in the $\delta_{1}$ vs $\delta_{2}$ plane, as defined in Eq. 11. Here we consider only SRI4 since it sets the strongest limit.
@@ 5 Numerical Analysis
$\delta_{1}(M_{N_{1}}\penalty 10000\ \rm{[GeV]})$  $10^{-3}$ , $0.5$] $([\approx M_{\chi},750])$
$\delta_{2}(M_{N_{2}}\penalty 10000\ \rm{[GeV]})$  $0^{-3} , 11$] $([\approx M_{N_{1}},1500])$
@@
In Fig. 13, we show the allowed range of $M_{N_{1}}$ vs $M_{N_{2}}$ with $c^{\prime}_{3}$ in the color pallet after imposing the relic density constraint as given by Eq. 14. We then impose the ATLAS mono-$\gamma$ limit as discussed in Sec. 4. Figure 13 shows that the mono-$\gamma$ search strongly constrain scenarios with large mass splitting between $M_{N_{1}}$ and $M_{N_{2}}$, these points are shown in red in the figure. Furthermore, as discussed above, the relic density constraint requires at least $N_{1}$ to be present in the early Universe, so that $\chi$ can be diluted through co-annihilation or inverse decay. Thus the mass difference between $M_{N_{1}}$ and $M_{\chi}$ cannot be too large, and we obtain an upper limit on $M_{N_{1}}<600\,\rm{GeV}$. It is important to notice that $N_{2}$ can also participate in DM $\chi$ dilution through co-annihilation or inverse decay. However since the contribution of $N_{1}$ is always present, the upper bound on $M_{N_{2}}$ is less stringent, specifically, $M_{N_{2}}\leq 1000\,\mathrm{GeV}$.
@@
In Fig. 13, we depict the allowed points after the relic density constraint and collider limit in the $\Omega_{\chi}h^{2}$ and $c^{\prime}_{3}$ plane with the color pallet corresponding to the DM mass $M_{\chi}$. The points overlapping with the red line satisfy the observed DM relic density within the $2\sigma$ range. As $c^{\prime}_{3}$ decreases, $N_{1,2}$ undergoes early chemical decoupling from the thermal bath, which in turn does not contribute in $\chi$ dilution neither through co-annihilation nor through co-scattering. We can see that for the range of $M_{\chi}$ that survived $c^{\prime}_{3}\leq 3\times 10^{-4}$ , i.e. $M_{\chi}>200\,\rm{GeV}$, corresponds to $M_{N_{1,2}}>M_{\chi}$. It happens because, for $M_{N_{1,2}}>200\,\rm{GeV}$ additional channels such as $N_{1}N_{2}\to t\bar{t}$ become kinematically accessible. In Fig. 13, we show the variation of the allowed points in $y^{\prime}_{11}$ - $y^{\prime}_{22}$ plane with the color pallet corresponding to $\Delta^{\chi}_{1}$. As evident from Fig. 13, for large $y^{\prime}_{11,22}$, $\chi$ thermalizes with the thermal bath due to rapid decay and inverse decay as $\Delta^{\chi}_{1}\to 0$, and the relic density is governed by the thermal decoupling of $N_{1,2}$. It is the region of parameter space where a co-annihilation process governs $\Omega_{\chi}h^{2}$. On the contrary, for $y^{\prime}_{11,22}<10^{-8}\,\rm{GeV^{-1}}$, $\chi$ undergoes early chemical decoupling and $\Omega_{\chi}h^{2}$ gets frozen when $\chi$ inverse decay to $N_{1,2}$ ceases. Consequently, in this region of parameter space, the value of $\Delta^{\chi}_{1}>0.5$, indicating that the DM relic density is predominantly determined by the co-scattering mechanism. In Fig 13, we show the discovery prospects via the mono-$\gamma$ signal at the HL-LHC. The dotted-color points in the $M_{N_{1}}$ and $M_{N_{2}}$ plane can be probed in HL-LHC while the red-crossed points remain out of HL-LHC reach. These points correspond to lower values of $M_{N_{2}}-M_{N_{1}}$ and/or lower value of $c^{\prime}_{3}$, which leads to suppressed signal efficiency and/or suppressed production cross section for the signal $pp\to N_{1}N_{2}\to\gamma+E_{T}^{\rm miss}$.
@@ 6 Summary and Conclusion
In this work, we investigate the dark matter production mechanism in an EFT framework where we extent the SM with two singlet fermions $N_{1,2}$ and a real singlet scalar $\chi$, which are $Z_{2}$ odd particles. The singlet scalar $\chi$ serves as a dark matter candidate. Both $\chi$ and $N_{1,2}$ are thermal particles and their interactions in the early Universe controls the DM abundance. We emphasize the role of dimension-5 operators in the co-scattering mechanism. The key parameters on which DM dynamics depend are the masses $M_{N_{1,2}},M_{\chi}$ and couplings $\lambda,c_{3}^{\prime},y^{\prime}_{11,22}$. Note that, we have assumed $c_{1,2}^{\prime}=0$ throughout our analysis.
@@
As concerns the collider constraints, we find that the dimension-5 operators lead to sizable production of $N_{1,2}$ at the LHC and that the most stringent constraint is coming from the mono-$\gamma$ signature via $pp\to N_{1}N_{2}(\to N_{1}\gamma)\to\gamma+E_{T}^{\rm miss}$. Following the ATLAS analysis, we calculate the sensitivity of this signal which primarily depends on the mass difference $M_{N_{2}}-M_{N_{1}}$ (or the photon energy) and $c_{3}^{\prime}$. We calculate the mono-$\gamma$ limit on $c_{3}^{\prime}$ in the $\delta_{2}-\delta_{1}$ plane, where $\delta_{1}$ dictates the relative mass splitting between $M_{N_{1}}$ and $M_{\chi}$. The stronger bound on $c_{3}^{\prime}$ is found for larger $\delta_{2}$ since the energy of the photon grows with $\delta_{2}$.